# Towards a universal QAOA protocol: Evidence of a scaling advantage in solving some combinatorial optimization problems

### v2

@@ II.3 COPs
A detailed description of some COPs used in this work can be found in the appendix of [18]. We pick 5 random instances for different problem sizes. For the TSP, we use instances with 3, 4, 5, and 6 cities (9, 16, 25, and 36 qubits), where the distances between cities are symmetric and randomly chosen from a uniform distribution with values between 0.1 to 1.1. In the BPP, we consider scenarios involving 3, 4, 5, and 6 items (12, 20, 30, and 42 qubits). The weight of each item is randomly chosen from 1 to 10, and 20 is the maximum weight of the bins. The W-MaxCut, 3-MaxCut, MIS, KP, and PO problem sizes are given by the number of qubits and chosen to be 4, 6, 8, 10, 12, 14, 16, 20, 25, 30, 35, and 40. For W-MaxCut and MIS we also include cases with 42 qubits.
For W-MaxCut problem simulations, we use randomly weighted edges with weights chosen uniformly between 0 and 1 and a probability of having an edge between any two vertices as 70%. One of these cases with its optimal solution is presented in Fig. 1-(d). To test the conjecture about the linear time complexity of LR-QAOA, we use fully connected random W-MaxCut, MaxCut, Max-3-SAT, and Max-2-SAT problems. The weights for W-MaxCut are chosen from a uniform discrete distribution from 0 to 1000 in steps of 1. For random MaxCut problems, i.e., with a constant edge weight of 1, we vary the probability of having edges between nodes from 5% to 95%. We simulate 10 random problems from problem sizes between 15 to 35 qubits. For Max-2-SAT, we use 20 random instances up to 20 qubits and 10 random cases of Max-3-SAT with problem sizes between 15 to 25 qubits. A description of the Max-3-SAT problem setup is left in Sec. A.1. For MIS, edges between nodes are randomly selected with a 40% probability of having an edge. For KP problems, item values range from 5 to 63, weights from 1 to 20, and maximum weight is set to half of the sum of all weights. Finally, for PO, correlation matrix values are chosen from $[-0.1,0,0.1,0.2]$, asset costs varying between 0.5 and 1.5, and the budget is set to half of the total assets cost.
For the inequality constraints in the KP, PO, and BPP, we use the unbalanced penalization approach [28, 30]. For the $probability(x^{*})$ using unbalanced penalization, our focus is on finding the ground state of the cost Hamiltonian, since we are interested in knowing the LR-QAOA performance in finding the ground state of the Hamiltonian and there is no guarantee that the optimal solution of the original problem is encoded in the ground state of the Hamiltonian (see also the discussion in [28]).
@@
where $n$ is the number of samples, $x_{i}$ the $ith$ bitstring obtained from LR-QAOA, and $C(x)$ is the cost function of W-MaxCut, $x^{*}$ is the optimal bitstring, $C(x^{*})$ is the maximum cut, $w_{kl}$ is the weight of the edge between nodes $k$ and $l$, and $x_{k}$ is the kth position of the $x$ bitstring.
@@
II.4 Classical solvers
We compare LR-QAOA, SA [32], and CPLEX’s spatial B&B [33] using TTS for the W-MaxCut problem. The TTS is given by
$TTS=\mathrm{T}\frac{\ln(1-p_{d})}{\ln(1-probability(x^{*}))},$ (12)
where $p_{d}$ is the target probability and $\mathrm{T}$ is the time complexity of the algorithm. For LR-QAOA, the time complexity comes from the number of layers used $\mathrm{T}_{LR-QAOA}=p$ while for SA it comes from the number of annealing steps and the number of transition states generated. We evaluate $N_{q}$ transition states at every annealing step and therefore the time complexity for SA is $\mathrm{T}_{SA}=SN_{q}$ where $S$ is the number of annealing steps. We consider two strategies one with the number of steps constant $S=5$ and another with the number of steps increasing linearly with the problem size $S=N_{q}$. For the case of B&B of CPLEX, for the sake of comparison, we consider the number of iterations to be the TTS.
II.5 Performance diagram
The performance diagram, introduced in [11], is a clear way to visualize the LR-QAOA performance. Figure 3-(a) shows the performance diagrams of a 10-qubit MIS problem. In the green-shaded ’annealing’ region, LR-QAOA behaves like a continuous QA protocol. This region is delimited by small $\Delta$ and large $p$ which guarantees yielding the optimal solution. The two points (black and white) marked in Fig. 3-(a) shows two LR-QAOA protocols with similar $probability(x^{*})$. The white point is an LR-QAOA with $p=300$ while the black point shows a LR-QAOA of $p=17$.
In Fig. 3-(b), we present the performance diagram of a 10-qubit W-MaxCut using LR-QAOA. The shaded red region in this plot illustrates the ’ridge’ as defined by [11]. The ridge region is characterized by consistently improving performance as the number of layers grows. The two points in this plot show similar performance, with the black point achieving a $probability(x^{*})>0.65$ with just $p=20$, while the white point in the annealing region requires a higher $p=300$. In appendix A.2, it is shown a 2-D version of the performance diagram for different problems.
Figure 3: Performance diagram using LR-QAOA on (a) a 10-qubit MIS and (b) a 10-qubit W-MaxCut problem. The y-axis represents $\Delta_{\gamma,\beta}$, and the x-axis is the number of LR-QAOA layers. The colors represent the $probability(x^{*})$ landscape, i.e., the probability of finding the optimal solution with darker red approaching the $100\%$. Circles in white and black represent similar $probability(x^{*})$ cases. The arrows connecting them are guiding elements.
II.6 Hamiltonian normalization
The Hamiltonian normalization is one important step in LR-QAOA. As we show, every eigenvalue rotates accordingly to Eq. 7, which means that the normalization limits the rotation angle fixing the ridge region to a specific location in the performance diagram. The general form of the COP’s Ising Hamiltonian is given by
@@ II.6 Hamiltonian normalization
Mitigation: Hamming distance 1
In [18], we introduce the mitigation technique used here. This involves applying a bitflip to each position within the output bitstring of samples from a quantum computer, to mitigate single-qubit bitflips. The computational overhead of this postprocessing method is $\mathrm{O}(NN_{q})$, where N represents the number of samples and $N_{q}$ is the number of qubits. The success of this method relies on the large probability LR-QAOA of amplifying solutions close to the optimal. The details of our proposed approach are described in Algorithm 1.
@@
II.7 Experimental details
@@
We perform different experiments to assess the practical performance of quantum technology to solve COPs using LR-QAOA. First, an experiment on ionq_aria for a 10-qubit W-MaxCut with 70% of random connections as described in Section II.3, this helps for the sake of comparison with a depolarizing noise model, Fig.  7. Additionally, different problems from 5 to 109 qubits were tested on ibm_brisbane using a W-MaxCut problem with a 1D-chain topology shown in Fig.  9-(a). We opt for a simple graph due to constraints posed by noise. Additionally, we provide an experimental comparison across three distinct IBM devices for a 109-qubit W-MaxCut problem. Finally, a comparison between ionq_aria, ibm_brisbane, and quantinuum_H2 is shown for a 25-qubit W-MaxCut problem, Fig. 9-(b).
The time of execution $t_{e}$ for 1D chain topology LR-QAOA protocol can be approximated to that of the 2-qubit gates. This is because single-qubit operations are a minority and their execution time is generally faster than 2-qubit gates. In ionq_aria the 2-qubit gates are executed sequentially so $t_{e}=t_{2q}N_{2q}p$ where $N_{2q}$ is the number of 2-qubit terms in the cost Hamiltonian. For the case of ibm_brisbane, the time of execution is $t_{e}=2t_{2q}p$. quantinuum_H2 can execute 4 2-qubit gates in parallel, hence, the execution time is $t_{e}=t_{2q}(N_{2q}/4)p$. The time per 2-qubit gate is $600\ \mu s$ on ionq_aria, and $660\ ns$ on ibm_brisbane. We could not find information about $t_{2q}$ for quantinuum_H2, but we assume it is similar to ionq_aria. Therefore, a 25-qubit W-MaxCut with 1D topology requires $14.4ms$, $3.6\ ms$, and $1.32\ \mu s$ for each layer using ionq_aria, quantinuum_H2, and ibm_brisbane, respectively. For each experiment on IBM devices, ionq_aria, and quantinuum_H2 we use 10000, 1000, and 50 samples, respectively.
II.8 Noise model
$\mathcal{E}[\rho]=(1-\lambda)\rho+\lambda\frac{I}{4},$ (14)
where $\lambda$ is the depolarizing error parameter, $I$ is a $4\times 4$ identity matrix, and $\rho$ is the density matrix of the 2-qubit system. In general, the action of a 2-qubit gate on a general density matrix can be expressed by
@@ III.1 Simulations
Figures 4-(a)-(d) show the probability of success for different COPs vs. the number of qubits. We test problems with 4 to 42 qubits using the LR-QAOA with $p=10$ to $400$, $\Delta_{\gamma}=0.6$, and $\Delta_{\beta}=0.3$. The error bars represent the minimum and maximum $probability(x^{*})$ over 5 random cases and the circles are the mean value. The colors represent the number of layers $p$ used. Figure 1-(d) shows one of the 5 random cases of W-MaxCut with 42-qubit of Fig. 4-(a). In this example just employing 50 layers of LR-QAOA results in an amplification of 12 orders of magnitude in the probability of success from approximately $2/2^{42}\approx 4.54\times 10^{-13}$ to $0.32$. To provide context, if we were to solve the same task using Grover’s algorithm [38], it would require approximately $\sqrt{2^{42}}/2\approx 2\times 10^{6}$ iterations of the oracle, along with the diffuser. The dashed lines in Figs. 4-(a)-(d) are added as guiding lines. They represent $1/2^{(\eta N_{q}/p)}$ where $\eta$ is a constant. The $\eta$ values used are (a) $\eta=2.8$, (b) $\eta=4$, (c) $\eta=2.5$, and (d) $\eta=3.5$. These results suggest that the $probability(x^{*})=1/2^{(\eta N_{q}/p)}$ for the random instances of different COPs analyzed. The implications are significant, for example, increasing the number of layers proportional to the number of qubits, i.e. $p=N_{q}$, one should observe that the probability of success must remain constant. We test if this is the case, in the Sec. III.2.
Figure 4-(e) shows the probability of success for different combinatorial optimization problems up to 42 qubits using the LR-QAOA protocol with $p=100$. The dashed line, $1/2^{N_{q}/2}$, is added as a guiding line, it reflects a quadratic speedup in the search space. The best performance is obtained for W-MaxCut with an average $probability(x^{*})=0.58$ at 40 qubits which compared to the optimal solution initial amplitude $2/2^{40}\approx 0.18\times 10^{-11}$ is an 11 orders of magnitude increment in only 100 layers of LR-QAOA. The lowest performance is for the TSP and BPP; for example, for a 6 cities problem (36 qubits) we get an average $probability(x^{*})=0.15$ which compared with the random guessing probability $1/2^{36}\approx 0.145\times 10^{-10}$ is a gain of 10 orders of magnitude. However, it is worth noting, that the number of feasible solutions to this problem is $(cities-1)!/2=60$, resulting in a random guessing probability within the feasible space of approximately $1/60\approx 0.0166$.
Figure 4-(f) shows the average performance of LR-QAOA for the W-MaxCut problem in terms of the fractional error $1-r$ vs. the number of qubits, where $r$ is the approximation ratio given by Eq. (10). The maximum number of qubits used is 40. Different colors represent the number of layers from 10 to 400. As expected, an increase in the number of layers leads to an increase in the average performance of LR-QAOA to find better solutions. Additionally, the relation of $1-r$ tends to stay constant with an increment in the number of qubits; this aligns with the minimum performance guarantee of MaxCut for a fixed number of QAOA layers [39].
@@
Figure 4: Probability of success for 5 random instances of (a) W-MaxCut, (b) MIS, (c) 3-MaxCut, and (d) PO. The error bars represent the minimum and maximum over the 5 cases. The different colors represent the number of LR-QAOA layers (see legend). Dashed lines represent the conjectured scaling $1/2^{(\eta N_{q}/p)}$ for each $p$. (e) Probability of success for 5 random instances for different COPs using a $p=100$ LR-QAOA. Markers represent the median value and error bars represent the Q1 and Q3 quartiles. The average performance of LR-QAOA as measured by (f) the fractional error, $1-r$, of the instances of W-MaxCut from (a), and (g) probability vs. cost of an instance of a 12-qubit PO solved using LR-QAOA with different $p$ from 10 to 200.
III.2 Conjecture
To add further evidence that $probability(x^{*})=1/2^{(\eta N_{q}/p)}$, we simulate LR-QAOA for fully connected W-MaxCut random problems for 100 cases and problem size ranging from 5 to 42 qubits using $p=N_{q}$. In Fig. 5, we show evidence supporting this hypothesis.
Figure 5-(a) shows the $probability(x^{*})$ for different problem sizes. The dots represent the probability of success for individual cases and the black circle is the mean value for a $\Delta_{\gamma,\beta}=0.5$. After reaching $N_{q}=25$, there is a decline in performance using $\Delta_{\gamma,\beta}=0.5$. To address this, we adjust $\Delta_{\gamma,\beta}$ to 0.4 for certain problem sizes, which leads to an improvement in performance. The data points corresponding to $\Delta_{\gamma,\beta}=0.4$ are shown with triangles. This suggests that one must reduce $\Delta_{\gamma,\beta}$ as the number of qubits increases. Interestingly, maintaining $N_{q}=p$ does not result in any significant performance loss. This observation strengthens the hypothesis that $probability(x^{*})=1/2^{(\eta N_{q}/p)}$. If this relation were to hold, it would mean that LR-QAOA is a polynomial time quantum algorithm for solving general instances of NP-Hard problems. However, it is still possible that there is a mild exponential decay for this problem that cannot be noticed at $\leq 42$ qubits. Even if this is the case, a scaling advantage is plausible.
Figure 5-(b) shows the approximation ratio for the same problems of (a). The decrement in performance is more notorious in terms of the approximation ratio for $\Delta_{\gamma,\beta}=0.5$. The modification of $\Delta_{\gamma,\beta}=0.4$ leads to a general improvement in terms of the quality of the solutions. It has been proven by Theorem 4.2 in [40], that it is NP-Hard to approximate MaxCut for a factor of $16/17\approx 0.941$. The dashed line represents $r=0.878$, the Goemans and Williamson (GW) algorithm [41], which is the best-known polynomial-time algorithm and according to the unique games conjecture also the best possible polynomial-time algorithm for this problem [42, 43].
Figure 5-(c) shows the $TTS_{99\%}$ for a $p_{d}=0.99$ (Eq. 12) of the W-MaxCut problems for SA and LR-QAOA. We include the CPLEX number of iterations needed to find the optimal solution for comparison. This plot shows the average TTS for the 100 W-MaxCut cases. We add three guiding lines of the perceived scaling of the 3 algorithms. CPLEX B&B seems to scale $\approx 2^{0.36N_{q}}$, SA $\approx 2^{0.19N_{q}}$, and LR-QAOA $\approx 2^{0.11N_{q}}$. For LR-QAOA, we also show an exponential growth line, but based on our conjecture, it could potentially grow linearly. Note that we do not tune the $\Delta_{\gamma,\beta}$ parameters in these cases, and even though the results are promising, further improvement can be achieved finding a function $\Delta_{\gamma,\beta}(Nq)$. In the case of LR-QAOA, at $N_{q}=42$, it has improved by two orders of magnitude the TTS compared to SA and B&B.
Fig. 5-(d) shows some instances of Max-2-SAT (triangles) and Max-3-SAT (circles) problems. The Max-3-SAT 10 random instances for problem sizes up to 25 qubits are chosen from a ratio of clauses/variables = 4.16 which is close to the critical region where the problems are known to be hard to solve [44, 45]. In this case, the relation $probability(x^{*})=1/2^{\eta N_{q}/p}$ seems to be valid on the average case. The Max-2-SAT problems used are obtained from a dataset of hard instances of this problem for SA and QA [46]. What makes these problems special is the number of degeneracies of the first excited state. For example, in the inset of this figure, it is shown a 20-qubit case. The optimal solution has 2 degeneracies while the first excited state has 6516. This causes SA and QA to get stuck in solutions close to optimal ones. In our case, LR-QAOA with $p=N_{q}$ does not hold a constant probability of success, so we can expect that $probability(x^{*})=1/2^{\eta N_{q}/p}$ relation does not hold always.
Fig. 5-(e) shows the LR-QAOA $p=N_{q}$ probability of success for 10 random cases of the MaxCut for up to 35 qubits. We use 7 different problems of MaxCut that we classify depending on the percentage of edges of the graph. The colors represent the average number of edges in the graph, meaning 0.05 having $5\%$ of all possible edges (dark blue) and $0.95$ a $95\%$ of all possible edges (dark red). In this case, on average the $p=N_{q}$ seems to hold $probability(x^{*})=1/2^{\eta N_{q}/p}$. However, the worst case highlighted with the red circle deviates considerably from the average performance. We explore this case and find the same characteristic that makes Max-2-SAT hard for LR-QAOA, a high degeneracy close to the optimal solution. Finally, Fig. 5-(f) shows the number of iterations needed to solve the same MaxCut problems of Fig. 5-(e) using CPLEX. In this case, it is classically hard to find the optimal solution for densely connected graphs with a worse-than-quadratic speedup for edge probabilities from $80\%$ to $95\%$.
Figure 5: Evidence for conjecutre 1. The problems shown here are solved using LR-QAOA with $p=N_{q}$ (a) Probability of success of fully connected W-MaxCut problem using 100 random W-MaxCut problems up to 40 qubits and 10 problems for 42 qubits. The different vertically stacked dots represent individual cases and the circles (triangles) indicate the mean value $\Delta_{\beta,\gamma}=0.5$ ($\Delta_{\beta,\gamma}=0.4$). The dotted line represents the probability of finding the optimal solution by random guessing and the dashed line represents the probability of finding the optimal solution if the search space is reduced by a quadratic speedup algorithm. (b) Approximation ratio for the W-MaxCut problems. The dashed line (GW) represents the approximation ratio using the best-known polynomial-time classical algorithm for the same problem. (c) TTS for SA and LR-QAOA compared to CPLEX’s number of iterations. Three guiding lines are added following the perceived scale of the different algorithms. (d) 20 hard cases of Max-2-SAT problems (triangles) for problem sizes up to 20 qubits and 10 random cases of Max-3-SAT problems (circles) for problems up to 25 qubits. The inset shows the degeneracies of the 4 first eigenvalues for the Max-2-SAT problem highlighted in red. (e) 10 random MaxCut instances for problem sizes up to 35 qubits. The different colors mean the average percentage of edges compared to a fully connected graph. Therefore, darker blue, $0.05$, means a $5\%$ of having edges. The circles represent the mean value and the dots represent the worst case. The inset shows the degeneracies of the solutions in the worst case. (f) The number of iterations needed to find the optimal solution using CPLEX for the MaxCut problems. The colors follow the same pattern as in (e).
With this extra information at hand, we present the following conjecture.
Conjecture 1.
For a given COP solved using LR-QAOA, there is a $\Delta_{\gamma,\beta}$ for which the probability of finding the optimal solution scales according to $probability(x^{*})=1/2^{(\eta N_{q}/p)}$ for a constant $\eta$, $p$ LR-QAOA layers, and $N_{q}$ qubits, iff there is no high concentration of solutions near the optimal solution.
This conjecture is helpful as a first attempt to know the expected performance of LR-QAOA for a given problem. It may suggest that LR-QAOA can find solutions to general instances of NP-Hard problems in polynomial time. However, the evidence involves solutions of a limited size and it is still possible that there is a mild exponential decay for problems like W-MaxCut. In the case of problems with suboptimal solutions densely concentrated close to the optimal, there is a noticeable exponential decay for $p=N_{q}$. In this scenario, for instance, in the Max-2-SAT problem, LR-QAOA still gives a quadratic speedup. In this case, the high degeneracy causes the dense concentration of solutions close to the optimal one. Similarly, in the TSP or BPP, the concentration of solutions occurs due to penalization terms concentrating the valid solutions space on a small region. We suspect that in the limit, i.e., a random oracle with only one ground state and the first excited state containing all other possible solutions, like the one in [47], one would need $p=2^{\eta N_{q}}$ layers to amplify the solution to some threshold for a constant $\eta$, similar to the performance guarantee in Grover’s algorithm.

### v3

@@ II.3 COPs
A detailed description of some COPs used in this work can be found in the appendix of [20], and for the Max-3-SAT is presented in the Supplementary Note 6. For them, we use a normalization technique described in Sec. II.4. We pick 5 random instances for different problem sizes. For the TSP, we use instances with 3, 4, 5, and 6 cities (9, 16, 25, and 36 qubits), where the distances between cities are symmetric and randomly chosen from a uniform distribution with values between 0.1 to 1.1. In the BPP, we consider scenarios involving 3, 4, 5, and 6 items (12, 20, 30, and 42 qubits). The weight of each item is randomly chosen from 1 to 10, and 20 is the maximum weight of the bins. The WMaxcut, 3-Maxcut, MIS, and PO problem sizes are given by the number of qubits and chosen to be 20, 25, 30, 35, and 40.
For WMaxcut problem simulations, we use randomly weighted edges with weights chosen uniformly between 0 and 1 and edge density, $E_{d}=0.7$. One of these cases with its optimal solution is presented in Fig. 1-(d). To test the scaling of LR-QAOA, we use a fully connected random WMaxcut with weights chosen from a uniform discrete distribution from 0 to 1000 in steps of 1. For MIS, edges between nodes are randomly selected with $E_{d}=0.4$. For KP problems, item values range from 5 to 63, weights from 1 to 20, and the maximum weight is set to half of the sum of all weights. Finally, for PO, correlation matrix values are chosen from [0.1,0,0.1,0.2][-0.1,0,0.1,0.2][ - 0.1 , 0 , 0.1 , 0.2 ], asset costs varying between 0.5 and 1.5, and the budget is set to half of the total asset cost.
For the inequality constraints in the KP, PO, and BPP, we use the unbalanced penalization approach [29, 31]. In this approach, two penalty terms in the QUBO are tuned following the characteristics of the inequality constraints and the objective function. Consequently, any variation in the parameter range necessitates a re-tuning of the penalty terms to maintain performance. For the $probability(x^{*})$ using unbalanced penalization, our focus is on finding the ground state of the cost Hamiltonian, since we are interested in knowing the LR-QAOA performance in finding the ground state of the Hamiltonian and there is no guarantee that the optimal solution of the original problem is encoded in the ground state of the Hamiltonian (see also the discussion in [29]).
From the problems tested, MIS, BPP, TSP, Maxcut, WMaxcut, KP, PO, and Max-3-SAT are NP-hard [2, 25], with varying structural properties and practical solution approaches. Some of them admit effective approximation schemes and are commonly addressed using heuristics or dynamic programming in restricted cases. In particular, MIS and PO have been included in a list of 10 classical hard problems that might benefit from quantum algorithms [32].
@@
where $n$ is the number of samples, $x_{i}$ the $ith$ bitstring obtained from LR-QAOA, and $C(x)$ is the cost function of WMaxcut, $x^{*}$ is the optimal bitstring, $C(x^{*})$ is the maximum cut, $w_{kl}$ is the weight of the edge between nodes $k$ and $l$, and $x_{k}$ is the kth position of the $x$ bitstring.
@@
II.4 Hamiltonian normalization
The Hamiltonian normalization is one important step in LR-QAOA. As we show, every eigenvalue rotates accordingly to Eq. 7, which means that the normalization limits the rotation angle, fixing the ridge region to a specific location in the performance diagram [11] (See Supplementary Note 7). The general form of the COP’s Ising Hamiltonian is given by
@@ II.4 Hamiltonian normalization
II.5 Classical solvers
To assess the performance of LR-QAOA, we compare its scalability to simulated annealing (SA) [34], TABU search [35], and CPLEX’s spatial B&B [36]. We selected TABU search because it has been shown to outperform other solvers in finding optimal solutions to Maxcut problems [37]. The improved performance in the TABU search can be attributed to a TABU list that prevents revisiting previous solutions and therefore mitigates local minimum problems. We use time-to-solution (TTS) as a metric to compare the resources needed to find the optimal solutions to fully connected WMaxcut. The $TTS$ is given by
$TTS_{p_{d}}=T\frac{\ln(1-p_{d})}{\ln(1-probability(x^{*}))},$ (13)
where T is the time needed to get one sample, $p_{d}=0.99$ is the target probability, i.e., the confidence level that the optimal solution is sampled at least once with 99% certainty. T in SA and TABU depend on the number of sweeps, with 1 sweep representing a full update cycle over all variables. In experiments, we vary the number of sweeps from 50 to 500.
For SA, we use the dwave-neal [38] and for TABU we use dwave-tabu [39], both performant C++-based libraries that use Python as an interface. In the case of the CPLEX solver, we use docplex [40] Python interface of CPLEX. All the algorithms run on a MacBook Pro equipped with an Apple M1 chip. The code used to run the given cases can be found at [41].
The case of LR-QAOA, the $T=t_{2q}(2N_{q}+2)p$ with $t_{2q}$ is the 2-qubit gate time, and the time to execute one layer of QAOA scales as $O(2N_{q}+2)$ based on a flexible scheme that can be run in a 1D chain of qubits [42]. The $t_{g}$ for most superconducting-based QPUs is on the order of nanoseconds.
II.6 Noise model
$\mathcal{E}[\rho]=(1-\lambda)\rho+\lambda\frac{I}{4},$ (14)
where $\lambda$ is the depolarizing error parameter, $I$ is a $\times 4$ 4 identity matrix, and $\rho$ is the density matrix of the 2-qubit system. In general, the action of a 2-qubit gate on a general density matrix can be expressed by
To test how noise affects the LR-QAOA solution for a given problem, we use the following relation,
$p_{ovl}=\frac{probability(x^{*})^{QPU}-probability(x^{*})^{r}}{probability(x^{*})-probability(x^{*})^{r}},$ (16)
where $p_{ovl}$ is the overlap between the ideal success probability $probability(x^{*})$ and the one obtained in the real device, $probability(x^{*})^{QPU}$, normalized by the random sampler success probability, $probability(x^{*})^{r}$. Additionally, we define the accumulated error in the circuit using
$\varepsilon_{acc}=N_{g}\lambda$ (17)
where $N_{g}$ is the number of 2-qubit gates involved in the circuit. Using this relation, we find that a model that describes $p_{ovl}$ is
$p_{ovl}=2^{-k_{0}\varepsilon_{acc}},$ (18)
where $k_{0}$ is a fitting parameter that depends on the problem. In results, we show that this approach can be applied to superconductive and trapped ion-based QPUs, obtaining a good match of experimental results for both devices.
II.7 Mitigation: Hamming distance 1
In [20], we introduce the mitigation technique used here. This involves applying a bitflip to each position within the output bitstring of samples from a quantum computer, to mitigate single-qubit bitflips errors. The computational overhead of this postprocessing method is $\mathrm{O}(NN_{q})$, where N represents the number of samples and $N_{q}$ is the number of qubits. While this mitigation technique can correct errors coming from the readout of the quantum device, it is also an optimization step that can completely obscure the optimization coming from the LR-QAOA algorithm. Therefore, it is important to compare the results against those obtained from a random sampler using the same mitigation technique, which is included in our main results. The details of our proposed approach are described in Algorithm 1.
@@ II.7 Mitigation: Hamming distance 1
II.8 Experimental details
We use random fully connected WMaxcut from 5 to 15 qubits. We run experiments on ibm_fez and H1-1E. For the case of ibm_fez, we use the parity twine chains (PTC) [44, 42] strategy to encode the LR-QAOA quantum circuit into a 1D-chain of qubits of the QPU. H1-1E is a 20-qubit emulator of the Quantinuum H1-1 QPU. In these experiments, $\Delta_{\gamma}=\Delta_{\beta}=0.6$, the number of samples is 1000.
@@ II.8 Experimental details
We perform different experiments to assess the practical performance of quantum technology to solve COPs using LR-QAOA. First, an experiment on ionq_aria for a 10-qubit WMaxcut with 70% of random connections as described in Section II.3, this helps for the sake of comparison with a depolarizing noise model. Additionally, different problems from 5 to 109 qubits were tested on ibm_brisbane using a WMaxcut problem with a 1D-chain topology shown in Fig.  8-(a). We opt for a simple graph due to constraints posed by noise. Additionally, we provide an experimental comparison across three distinct IBM devices for a 109-qubit WMaxcut problem. Finally, a comparison between ionq_aria, ibm_brisbane, and quantinuum_H2 is shown for a 25-qubit WMaxcut problem, Fig. 8-(b).
[…]
@@ III.1 Simulations
Figure 3: Probability of success for 100 random instances of (a) WMaxcut, (b) MIS, (c) 3-Maxcut, and (d) PO. The shaded region represents the quartiles 1 and 3 over the 100 cases. The different colors represent the number of LR-QAOA layers (see legend). Dashed lines represent the conjectured scaling $2^{-\eta(p)N_{q}+C}$ for each $p$. (e) Fitted $\eta$ vs. number of LR-QAOA layers for the mean value of the problems. (f) Relative error of $probability(x^{*})$ calculated using the fitting parameters vs. the number of LR-QAOA layers.
Figures 3-(a)-(d) show the average probability of success for different COPs vs. the number of qubits. We test 100 random problems with 10 to 35 qubits using the LR-QAOA with $p=10$ to 200200200. The $\Delta_{\gamma}$ and $\Delta_{\beta}$ values are scanned for each problem size for one instance, and that value is used for the remaining 99 instances. Although this methodology is not optimal, since ideally, individual parameters should be tuned for each case, we adopt this approach to manage the growing simulation costs associated with a large number of qubits. Moreover, this strategy highlights an important feature of LR-QAOA, even without fine-tuning parameters for each instance, the algorithm consistently achieves good performance. In Supplementary Note 1, we present in detail the methodology and the values used. In these problems, we observe that the number of layers affects the scale of the probability of success with a relation that can be described by
$probability(x^{*})=2^{-\eta(p)N_{q}+C},$ (19)
with an $\eta(p)$ that is a function of $p$, and $C$ a constant. The perceived scaling still needs to be corroborated at a larger problem size to confirm that the probability of success indeed decreases exponentially. A similar scaling QAOA behaviour has been observed for k-SAT problems in [27]. If this holds, it means that there is an exponential improvement in LR-QAOA achieved by increasing the number of layers linearly. This does not necessarily mean that the problems are hard and that the best classical solvers for them exhibit the perceived exponential scaling of LR-QAOA.
Figure 3-(e) shows the fitting values of $\eta(p)$ for each $p$ using the information of Figs 3-(a)-(d). The four models exhibit similar $\eta(p)$ behavior, decaying quickly with the number of layers. The fitted $\eta(p)$ implies an exponential improvement as the number of layers increases. For example, if the WMaxcut scale holds at $N_{q}=100$, using LR-QAOA with $p=10$ the $probability(x^{*})=2\times 10^{-7}$ while $probability(x^{*})=0.2$ with $p=200$. Figure 3-(f) shows the relative error of the difference between the value predicted by the fitting curve and the actual values for each $p$, i.e., $\varepsilon=|1-\frac{2^{-\eta(p)N_{q}+C}}{probability(x^{*})}|$. As the number of layers increases, the error decreases in 3 out of 4 cases and remains below 6%.
In Supplementary Note 2, we present the details about the scaling factor $\eta(p)$ and its relation with the minimum number of qubits used in the cutoff of the fitting function. The previous version of this paper included a conjecture that the probability of success scales as $probability(x^{*})=2^{-\eta N_{q}/p}$, with the further evidence of this section, the model of the scale of $probability(x^{*})$ has changed. Previous results and extended simulations are presented in Supplementary Note 3. In the next section, we present a comparative analysis of LR-QAOA and several classical solvers.
III.2 Scaling comparison
Figure 4: Scaling comparison in terms of TTS vs. number of qubits for classical solvers and LR-QAOA for fully connected random problems of WMaxcut. (a) The left plot shows the TTS of SA vs. TABU search for different numbers of qubits. Each marker represents a problem. The right plot shows the TTS vs. the 100 random cases sorted by TTS of SA for 40-qubit size problems. The shaded region highlights the 20 problems with the longest TTS. (b) TTS of SA, TABU search, and CPLEX. Three configurations are used for SA and TABU in each case: 50, 100, 200, and 500 sweeps, and the lowest TTS is used for the scaling. (c) TTS of LR-QAOA for $p=50$ and 100100100, the scaling is taken over the minimum TTS at each time-point. The markers represent the median value, and the error bars represent the first and third quartiles of the 20 cases distribution.
In Figure 4(a)-left, we show the TTS of SA vs. TTS of TABU search. There is a high correlation coefficient (PCC=0.7) [48] between the solvers’ TTS, indicating that random problems requiring longer TTS for one solver tend to require longer TTS for the other as well. In contrast, the PCC between TABU and CPLEX is 0.23, reflecting a weak correlation; thus, what is considered difficult for TABU is not necessarily difficult for CPLEX. Based on this information, we select the 20 cases with the longest TTS for SA out of the 100 random instances generated for each problem size. The shaded region in Fig. 4(a)-right highlights the problems selected for the 40-qubit problem size.
Figure 4(b) shows the scaling of SA and TABU search, and CPLEX B&B in seconds. We use sweeps ranging from 50 to 500 for the heuristic solvers and choose the minimum TTS in each problem size case. In the case of SA, the best TTS is found using 50 to 100 sweeps. This is a consequence of the time required to implement the sweeps. While more sweeps generally lead to better solutions, the improvement does not always justify the additional evaluation time. Therefore, there is a tradeoff between the number of sweeps and the evaluation time, indicating that an optimal number of sweeps exists for a given problem size.
The case of TABU is similar to SA; for a small number of qubits, the optimal number of sweeps is around 100. However, as the number of qubits increases, the best configurations shift. At $N_{q}=40$, both 200 and 500 iterations yield similar TTS. The scaling for SA is slightly better than TABU search, with a shorter TTS in all cases. At the problem sizes considered, the advantage of TABU search in maintaining a list of previously visited solutions does not appear to be necessary. As a result, the additional computational cost of comparing against this list at each iteration might affect the TTS. The TTS of CPLEX is several orders of magnitude higher than that of the other solvers, and the corresponding fitting function does not appear to be reliable.
Figure 4(c) presents the TTS of LR-QAOA in arbitrary units, which must be rescaled according to the two-qubit gate time, $t_{g}$, of a given quantum computer. We use $p=50$ and $p=100$ and choose the best TTS from them. For visualization, we use a $t_{g}=2.5\times 10^{-9}$ that matches the time of TABU search at $N_{q}=20$ and corresponds to a gate time of $t_{g}=2.5\ ns$. Comparing the models, LR-QAOA shows a potential scaling advantage over the other solvers. Achieving competitive scaling with LR-QAOA would likely require depths beyond $p>100$. The relative error of the perceived scaling is 0.211 for CPLEX, 0.023 for SA, 0.019 for TABU, and 0.011 for LR-QAOA.